4- Imagine the situation which a vehicle pulls out onto a single-lane highway that has a flow rate of 400 veh/h (Poisson distributed). The driver of the vehicle does not look for oncoming traffic. Road conditions and vehicle speeds on the highway are such that it takes 2.2 seconds for an oncoming vehicle to stop once the brakes are applied. Assuming a standard driver reaction time of 2.5 seconds,
a) (8%) what is the probability that the vehicle pulling out will get in an accident with oncoming traffic?
b) (10%) What should be the reaction and braking times of oncoming vehicles so that the probability of an accident is equal to 0.25?
Answer #1

If the vehicle headways between successive vehicles are greater than (2.5+2.2) =4.7sec then the driver pulling out will not b

b) Given Prch <t)=0.25 Pr(h<t)=1-eit 0.25=1-8-0. e-0.1111xt = 0.75 -0.111 lxt t = 2.589 sec Therefore Reaction time is Reacti

Hope this helps

Please feel free to comment on any further queries

All the best :)